# Point configurations in sets of sufficient topological structure and a topological Erdős similarity conjecture

### v1

@@ Appendix A Examples and properties of Baire sets
    $A=G\cup P$ for some $G_{\delta}$ set $G$ and some meager set $P$.
@@
    $A=F\setminus P$ for some $F_{\sigma}$ set $F$ and some meager set $P$.
@@
The most well-known example of a non-measurable set is also an example of a non-Baire set. It also shows the necessity of the assumption in our results that our sets are Baire.
Example A.5 (A non-meager set for which the conclusion of Theorem 2.1 fails).
@@
would be meager (this is pointed out in [20], where it is used to prove directly that $A$ is not Baire). Let $F=\{x,y\}\subset\mathbb{R}$, where $x-y\in\mathbb{Q}$. By definition, for any $t\in\mathbb{Q}$ and $z\in\mathbb{R}$ we have $(tx+z)\sim(ty+z)$. By construction, $A$ cannot contain two distinct equivalent numbers. It follows that $\Delta_{F}(A)$ does not contain any rational number, and hence cannot have nonempty interior (in light of Theorem 2.1, this gives a less-direct proof that $A$ is not Baire).
There is a broad literature on fractal sets which avoid given configurations, and generally, such a set must be meager. This makes it very easy to import examples which show that Theorem 2.1 (and hence the more general Theorem 3.3) fail if a set is assumed to be Baire but not necessarily non-meager. For example, for any dimension $d$, there are well-known constructions of compact sets $A$ (of full dimension) such that $A-A$ has empty interior. In the language of this paper, this is the same as saying $\Delta_{F}(A)$ has empty interior for any set $F$ with at least two points. Beyond this, we can give examples of $3$-point sets $F$ and Baire sets $A$ of positive Hausdorff dimension where the set $\Delta_{F}(A)$ is actually empty (rather than simply having empty interior). To do this, we use a construction of Mathe.
@@
Example A.7 (An meager set with positive Hausdorff dimension for which the conclusion of Theorem 2.1 fails).
@@
$a_{n}|y-x|^{2m}+a_{n-1}|y-x|^{2n-2}|z-x|^{2}+\cdots+a_{1}|y-x|^{2}|z-x|^{2n-2}+a_{0}|z-x|^{2n}=0.$
Thus, the points of $F$ satisfy a polynomial equation with rational coefficients. Moreover, by homogeneity and translation invariance, one can easily verify that the equation remains satisfied if $x,y,z$ are replaced by $tx+w,ty+w,tz+w$, respectively, for any $t>0$ and $w\in\mathbb{R}^{d}$. By Theorem A.6, it follows that there exists a compact set $A\subset\mathbb{R}^{d}$ such that $\Delta_{F}(A)=\varnothing$. Such a set $A$ must have empty interior, so it is nowhere dense and hence meager. However, because it is compact, it is still a Baire set by Corollary A.2. This shows the non-meager assumption in our theorems is necessary.
Our next example shows that the assumption that $F$ is countable in Theorem 2.1 is necessary.
@@
As metioned in the introduction, Gallagher, Lai, and Weber [4, Theorem 1.5] showed that Cantor sets on $\mathbb{R}$ with positive Newhouse thickness are not universal in the set of dense $G_{\delta}$ sets. In particular, they showed that there exists a dense $G_{\delta}$ set $G$ of full measure such that, for all Cantor sets $C$ with positive Newhouse thickness, $G$ does not contain an affine copy of $C$. The classic middle-thirds Cantor set, for example, has positive Newhouse thickness, and hence is not universal in the set of dense $G_{\delta}$ sets. Note that their $G$ is Baire, since it is a $G_{\delta}$ set, and $G$ is nonmeager since every dense $G_{\delta}$ set is comeager and hence non-meager.
Now, we consider an example that shows that our main Euclidean result, Theorem 2.1, is immediate for a special class of non-meager sets, mainly co-meager sets.
Example A.9 (countable sets are universal in co-meager sets).
Suppose that $G$ is meager and $F$ is countable. Observe that the Minkowski sum $G+F$ is meager as this set can be expressed as the countable union of shifts of the meager set $G$. Since $\mathbb{R}$ is not meager (by the Baire category theorem), then $\mathbb{R}\neq G+F$. Similarly, $\mathbb{R}\neq G-F$. This is equivalent to $(x+F)\cap G=\emptyset$ for some $x\in\mathbb{R}$. It follows that, if $F$ is countable and $C$ is co-meager, then $F+x\subset C$ for some $x\in\mathbb{R}$.
@@
Example A.10 (A meager set with full measure; a second category set with zero measure).
Enumerate the rationals by $q_{1},q_{2},\cdots$, and let $O(i,j)$ denote the open interval with center $q_{j}$ and length $\frac{1}{2^{i+j}}$. Set $U_{i}=\bigcup_{j=1}^{\infty}O(i,j)$, and set $D=\bigcap_{i=1}^{\infty}U_{i}$. One may verify that $D$ is a set of measure zero, while $\𝐷M=\mathbb{R}\backslash D$ is meager (as the complement of an intersection of open dense sets; each $U_{i}$ contains $\mathbb{Q}$). In conclusion, $\mathbb{R}=M\bigcup D$ can be expressed as the disjoint union of a meager set with infinite measure and a co-meager set (second category) with zero measure.

### v2

@@ Appendix A Examples and properties of Baire sets
    $A=G\cup M$ for some $G_{\delta}$ set $G$ and some meager set $M$.
@@
    $A=F\setminus M$ for some $F_{\sigma}$ set $F$ and some meager set $M$.
@@
The most well-known example of a non-measurable set is also an example of a non-Baire set. It also shows the necessity of the assumption that $A$ Baire in Theorem 2.1.
Example A.5 (A non-meager set $A$ for which the conclusion of Theorem 2.1 fails).
@@
would be meager (this is pointed out in [23], where it is used to prove directly that $A$ is not Baire). Let $P=\{x,y\}\subset\mathbb{R}$, where $x-y\in\mathbb{Q}$. By definition, for any $t\in\mathbb{Q}$ and $z\in\mathbb{R}$ we have $(tx+z)\sim(ty+z)$. By construction, $A$ cannot contain two distinct equivalent numbers. It follows that $\Delta_{P}(A)$ does not contain any rational number, and hence cannot have nonempty interior (in light of Theorem 2.1, this gives an alternative proof that $A$ is not Baire). The reader who is interested in set theory should note that the Axiom of Choice is crucial in this construction. Indeed, there is a model of ZF without choice in which all sets are both measurable and Baire (see, for example, [11, Theorem 26.14])
Next, we consider a simple consequence of the Baire category theorem showing that countable sets are universal in the collection of dense $G_{\delta}$ sets. Since every dense $G_{\delta}$ set is comeager, and since every comeager set contains a dense $G_{\delta}$ set, we see that universality in the collection of dense $G_{\delta}$ sets is equivalent to universality in the collection of comeager sets.
Example A.6 (finite and countably infinite sets are universal in the collection of comeager sets and in the collection of full measure sets ).
[…]
A very similar argument shows that countable sets are full measure universal. If $P$ is countable and $A$ is a set of full measure, then $\𝐴P+\mathbb{R}\backslash A$ is countable so that $\𝐴P+\mathbb{R}\backslash A\neq\mathbb{R}$, and so there is a $t\in\mathbb{R}$ so that $P+t\subset A$.
There is a broad literature on fractal sets which avoid given configurations, and generally, such a set must be meager. This makes it possible to import examples which show that Theorem 2.1 (and hence the more general Theorem 3.3) fail if a set is assumed to be Baire but not necessarily non-meager. For example, for any dimension $d$, there are well-known constructions of compact sets $A$ of dimension arbitrarily close to $d$ such that $A-A$ has empty interior (see for example [3, Theorem 8.15]). In the language of this paper, this is the same as saying $\Delta_{P}(A)$ has empty interior for any set $P$ with at least two points. Beyond this, we can give examples of $3$-point sets $P$ and Baire sets $A$ of positive Hausdorff dimension where the set $\Delta_{P}(A)$ is actually empty (rather than simply having empty interior). To do this, we use a construction of Máthé.
@@
The following demonstrates the necessity of the assumption that $A$ is non-meager in Theorem 2.1. Further, it demonstrates that even if $A$ has positive Hausdorff dimension, the conclusion of this theorem need not hold.
Example A.8 (A meager set with positive Hausdorff dimension for which the conclusion of Theorem 2.1 fails).
@@
$a_{n}|y-x|^{2n}+a_{n-1}|y-x|^{2n-2}|z-x|^{2}+\cdots+a_{1}|y-x|^{2}|z-x|^{2n-2}+a_{0}|z-x|^{2n}=0.$
Thus, the points of $P$ satisfy a polynomial equation with rational coefficients. Moreover, by homogeneity and translation invariance, one can easily verify that the equation remains satisfied if $x,y,z$ are replaced by $tx+w,ty+w,tz+w$, respectively, for any $t>0$ and $w\in\mathbb{R}^{d}$. By Theorem A.7, it follows that there exists a compact set $A\subset\mathbb{R}^{d}$ such that $\Delta_{P}(A)=\varnothing$. Such a set $A$ must have empty interior, so it is nowhere dense and hence meager. However, because it is compact, it is still a Baire set by Corollary A.2. This shows the non-meager assumption in our theorems is necessary.
Our next example shows that the assumption that $P$ is countable in Theorem 2.1 is necessary.
@@
As mentioned in the introduction, Gallagher, Lai, and Weber [5, Theorem 1.4] proved that Cantor sets in $\mathbb{R}^{d}$ are not universal in the collection of dense $G_{\delta}$ sets (equivalently, comeager sets). In particular, this implies that Cantor sets are not universal in the collection of non-meager Baire sets.
@@
Example A.10 (A meager set with full measure and a second category set with zero measure).
Enumerate the rationals by $q_{1},q_{2},\cdots$, and let $O(i,j)$ denote the open interval with center $q_{j}$ and length $\frac{1}{2^{i+j}}$. Set $U_{i}=\bigcup_{j=1}^{\infty}O(i,j)$, and set $D=\bigcap_{i=1}^{\infty}U_{i}$. One may verify that $D$ is a set of measure zero, while $\𝐷M=\mathbb{R}\backslash D$ is meager (as the complement of an intersection of open dense sets; each $U_{i}$ contains $\mathbb{Q}$). In conclusion, $\mathbb{R}=M\bigcup D$ can be expressed as the disjoint union of a meager set with infinite measure and a comeager set (second category) with zero measure.